helo find local max and min i found partial derivations but having trouble finishing

Find the local maximum and minimum values f(x,y)= x3 - 3x + y2
Answer #1

Solution:- Given function is frany) = 33-38 +y2 - d (83_-37+ 42 ) = 342 -3 - Ja I d ra3–3a + y2)=ay - d ( 332-3) = 60 (99 )=0LUVU slove the equation © we get yao slove the equation ©, we get a = / (10) pairs | (-1,0) 1 stationary , 8=60 o ) S = 0 e tvalue of the given the local manimum function is a - value of the given 18 = -2 the local minimum